# Unified Performance Control for Non-Square Nonlinear Systems with Relaxed Controllability

### v1

@@ IV Stability Analysis
Now we give the stability analysis of the control laws designed in Section III-D for system (1).
@@
Consider the uncertain non-square MIMO nonlinear strict-feedback system (1) with actuator faults. Under Assumptions 1-4, if the control law (46) with the virtual controllers (21), (32) and (38) is applied, then the control objectives mentioned in Section II can be achieved.
@@
$V_{N}=V_{N1}+\frac{1}{2\lambda_{N}\sigma_{N}}{\tilde{\theta}_{N}}^{2},$ (48)
@@
$\bm{\mathcal{G}}_{N}\bm{A}^{T}=\,\frac{\bm{G}_{N}}{2}+\frac{\left(\bm{\mathcal{G}}_{N}\bm{A}^{T}-\bm{A}\bm{\mathcal{G}}_{N}^{T}\right)}{2},$ (50)
@@
where $\Upsilon=\min\{\frac{2\lambda_{\textrm{min}}(\bm{W}^{T}\bm{W})\kappa_{1}\lambda_{1}}{\lambda_{\textrm{max}}(\bm{P}_{1})},\frac{2\kappa_{k}\lambda_{k}}{\lambda_{\textrm{max}}(\bm{P}_{N})},\mu_{k}\}>0$ and $C=\Delta_{N-1}+\frac{\mu_{N}}{2\sigma_{N}}{\theta}_{N}^{2}+\frac{3}{2}>0$.
@@
Now we first prove that all signals in the closed-loop systems are bounded. It is seen from (52) that $V_{N}(t)\leq V_{N}(0)+\frac{C}{\Upsilon}=\Xi$, it follows that $V_{N}(t)$ is bounded for $t\in[0,\tau_{\textrm{max}})$. According to the definition of $V_{N}(t)$, it follows that ${\bm{\varepsilon}}_{i}$ and $\tilde{\theta}_{i}$ are bounded and satisfy ${\bm{\varepsilon}}_{1}\leq\sqrt{\frac{2\lambda_{\textrm{min}}(\bm{W}^{T}\bm{W})\Xi}{\lambda_{\textrm{max}}(\bm{P}_{1})}}$, ${\bm{\varepsilon}}_{k}\leq\sqrt{\frac{2\Xi}{\lambda_{\textrm{max}}(\bm{P}_{k})}}$ ($k=2,\ldots,N$), and $\tilde{\theta}_{i}\leq\sqrt{2\sigma_{i}\Xi}$ for $t\in[0,\tau_{\textrm{max}})$ with ${\lambda_{\textrm{max}}(\bm{P}_{k})}>0$, which further indicates that the parameter estimate $\hat{\theta}_{i}$ is bounded over the interval $t\in[0,\tau_{\textrm{max}})$. Since ${\bm{\varepsilon}}_{1}$ is bounded, and using the properties of $s_{j}$, it is shown that $-1\leq-\underline{\delta}_{j}<-\underline{\delta}_{1j}\leq\zeta_{j}(t)\leq\bar{\delta}_{1j}<\bar{\delta}_{j}\leq 1$ for $t\in[0,\tau_{\textrm{max}})$. As $\zeta_{j}=\frac{\eta_{j}}{\varphi_{j}}$ and $0<\varphi_{jf}<\varphi_{j}(t)\leq\varphi_{j0}\leq 1$, it is proved that there exist some constants $\underline{\eta}_{j}$ and $\bar{\eta}_{j}$ so that $-1<\underline{\eta}_{j}\leq\eta_{j}\leq\bar{\eta}_{j}<1$, which further implies from $e_{j}=\frac{l_{j}\eta_{j}}{\sqrt{1-\eta_{j}^{2}}}$ that the tracking error $e_{j}$ is bounded for $t\in[0,\tau_{\textrm{max}})$. As $e_{j}=x_{1j}-y_{dj}$, one has $x_{1j}$ is bounded since $y_{dj}$ is bounded for $t\in[0,\tau_{\textrm{max}})$. Hence, it is trivial to prove that $0<\underline{\mu}_{j}\leq\mu_{j}\leq\bar{\mu}_{j}$ and $0<\underline{r}_{j}\leq r_{j}\leq\bar{r}_{j}$, $0<\underline{w}_{j}\leq w_{j}\leq\bar{w}_{j}$ for $t\in[0,\tau_{\textrm{max}})$, with $\underline{\mu}_{j}$, $\bar{\mu}_{j}$, $\underline{r}_{j}$, $\bar{r}_{j}$,$\underline{w}_{j}$, $\bar{w}_{j}$ being some positive constants, which indicates that $\bm{W}$ and $\bm{V}$ are bounded. Since $x_{1}$ is bounded, it is shown from Assumption 1 and 4 that $\Phi_{1}$ is bounded, and thus $\bm{a}_{1}$ and $\dot{\hat{\theta}}_{1}$ are bounded by (21) and (22). Following this line of argument, $\bm{a}_{i}$ ($i=2,\ldots,N-1$), $\bm{u}$, $\dot{\hat{\theta}}_{k}$ ($k=2,\ldots,N$) and $\bm{x}_{k}$ are bounded over the interval $t\in[0,\tau_{\textrm{max}})$. Note that the boundedness of all signals relies on the design parameters, yet is independent of $\tau_{\textrm{max}}$, which reveals that $\tau_{\textrm{max}}$ can be extended to $\tau_{\textrm{max}}=\infty$. Thus, all signals in the closed-loop systems are bounded on $t\in[0,\infty)$.
@@
Next we prove that the tracking error is always within the predefined region for $t\in[0,\infty)$. As $s_{j}$ is bounded for $t\in[0,\infty)$, according to Lemma 1, it is ensured that $\mathcal{H}(-\underline{\delta}_{j}\varphi_{j}(t))<e_{j}(t)<\mathcal{H}(\bar{\delta}_{j}\varphi_{j}(t))$, $j=1,2,\ldots,m$. The proof is completed. $\square$
@@
It can be seen that the system (1) with output constraints is transformed to an unconstrained system (14) by introducing a unified prescribed performance function $\mathcal{H}$. Then, by applying the proposed virtual controllers (21), (32), (38) and actual controller (46) to ensure the stability of the transformed system, the predefined performance constraint on the tracking error is indirectly implemented. It is worth noting that by adjusting the key design parameters, the global performance and semi-global yet asymmetric performance can be achieved without changing the control structure. Furthermore, the unexpected actuator faults can be compensated automatically without resorting to any additional fault detection and diagnosis module.
@@
The controllability condition and controller constructed in this paper can be flexibly applied to a class of square and non-square systems. Specifically, square and non-square systems can be handled by simply choosing $\bm{A}=\bm{I}_{n}$ and $\bm{A}=[\bm{I}_{n},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{(m-n)}]$, respectively. Moreover, since the auxiliary matrix $\bm{P}_{i}$ is adaptively compensated and does not use in the controller, even if $\bm{g}_{i}$ satisfies the original strong controllability condition in [9], it is sufficient to compensate $\bm{P}_{i}$ as a identity matrix $\bm{I}_{n}$ without any additional operation.
@@ V Illustrative Example
Figure 4: The evolution of $e_{j}$ under $\underline{\delta}_{j}=1,\ \bar{\delta}_{j}=0.7$, $i=1,2,3$.
@@
$\bm{J}_{0}=\begin{bmatrix}40&1.2&0.9\\ 1.2&17&1.4\\ 0.9&1.4&15\end{bmatrix}\textrm{kg}\cdot\textrm{m}^{2},$
@@
$\bm{\rho}(t)=\begin{cases}\textmd{diag}\big{\{}0.9+0.1\sin(t),1-0.2\tanh(t),\\ \ 0.9+0.1\cos(t),0.8-0.2\sin(t)\big{\}},\ \ \,t\in(0,5]\\ \textmd{diag}\big{\{}0.8+0.2\sin(t),0.8-0.2\tanh(t),\\ \ 0.8+0.2\cos(t),0.8-0.1\tanh(t)\big{\}},\ t\in(5,\infty)\end{cases}$
$\bm{\upsilon}(t)=[0.01\tanh(2t),0.01\cos(t),0.01\sin(3t),0.01\cos(2t)]^{T}.$
The key design parameters are selected as $\underline{\delta}_{j}=\bar{\delta}_{j}=1$, $j=1,2,3$. The control parameters are given as: $\kappa_{1}=1$, $\sigma_{1}=0.01$ and $\mu_{1}=0.1$. The matrix $\bm{A}$ is setting as
$\bm{A}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&1\end{bmatrix}.$
The rate function is chosen as $\beta(t)=\exp(-0.9t)\cos^{2}(t)$, and $\varphi_{j0}=1$, $\varphi_{jf}=0.1$ and $l=0.9936$, $j=1,2,3$. The initial conditions are set as: $\bm{x}_{1}(0)=[-0.5,-0.5,-0.5]^{T}$(rad) and $[1,0.5,0.5]^{T}$(rad), $\hat{\theta}_{1}=0$. The simulation results are given in Figs.1 and 2. Fig.1 shows that the output tracking errors $e_{j}$, $j=1,2,3$, evolve within the prescribed performance bounds ($\mathcal{H}(-\varphi_{j}(t)),\mathcal{H}(\varphi_{j}(t))$). Fig.2 shows the boundedness of the control signal $u$.
To verify that the proposed control is able to achieve the asymmetric performance behaviors, we select the key design parameters as $\underline{\delta}_{j}=0.8,\ \bar{\delta}_{j}=1$ and $\underline{\delta}_{j}=1,\ \bar{\delta}_{j}=0.7$, $j=1,2,3$. The initial conditions are set as: $\bm{x}_{1}(0)=[-0.5,-0.5,-0.5]^{T}$(rad) and $[1,0.5,0.5]^{T}$(rad). Other parameters and conditions are kept consistent with the previous simulations. The responses of tracking errors are depicted in Fig.3 and Fig.4, which shows that the tracking errors are always evolve within the prescribed performance bounds, in line with the theoretical analysis.

### v2

@@ IV Stability Analysis
Now we give the stability analysis of the control laws designed in Section III-C for system (1).
@@
Consider the uncertain non-square MIMO nonlinear strict-feedback system (1) subject to actuator faults. Under Assumptions 14, if the control law (46) is implemented together with the virtual controllers given in (21), (32), and (38), then all signals in the closed-loop system remain bounded despite the presence of actuator faults, and the tracking error $\bm{e}_{j}$ is guaranteed to stay within the prescribed performance bounds ($\mathcal{H}(-\varphi_{j}(t)),\mathcal{H}(\varphi_{j}(t))$).
@@
$V_{N}=V_{N1}+\frac{1}{2\lambda_{N}\sigma_{N}}{\tilde{\theta}_{N}}^{2}$ (48)
@@
$\bm{\mathcal{G}}_{N}\bm{A}^{T}=\,\frac{\bm{G}_{N}}{2}+\frac{\left(\bm{\mathcal{G}}_{N}\bm{A}^{T}-\bm{A}\bm{\mathcal{G}}_{N}^{T}\right)}{2}$ (50)
@@
where $\Upsilon=\min\{\frac{2\lambda_{\textrm{min}}(\bm{W}^{T}\bm{W})\kappa_{1}\lambda_{1}}{\lambda_{\textrm{max}}(\bm{P}_{1})},\frac{2\kappa_{k}\lambda_{k}}{\lambda_{\textrm{max}}(\bm{P}_{N})},\frac{\mu_{i}}{\lambda_{i}}\}>0$ for $k=2,\ldots,N$, $i=1,\ldots,N$, and $C=\Delta_{N-1}+\frac{\mu_{N}}{2\sigma_{N}}{\theta}_{N}^{2}+\frac{3}{2}>0$.
@@
We first establish the boundedness of all signals in the closed-loop system. From (52), we have $V_{N}(t)\leq V_{N}(0)+\frac{C}{\Upsilon}=\Xi$, which implies that $V_{N}(t)$ remains bounded for all $t\in[0,\tau_{\textrm{max}})$. By the definition of $V_{N}(t)$, it follows that both ${\bm{\varepsilon}}_{i}$ and $\tilde{\theta}_{i}$ are bounded and satisfy ${\bm{\varepsilon}}_{1}\leq\sqrt{\frac{2\lambda_{\textrm{min}}(\bm{W}^{T}\bm{W})\Xi}{\lambda_{\textrm{max}}(\bm{P}_{1})}}$, ${\bm{\varepsilon}}_{k}\leq\sqrt{\frac{2\Xi}{\lambda_{\textrm{max}}(\bm{P}_{k})}}$ ($k=2,\ldots,N$), and $\tilde{\theta}_{i}\leq\sqrt{2\sigma_{i}\Xi}$ for $t\in[0,\tau_{\textrm{max}})$ with ${\lambda_{\textrm{max}}(\bm{P}_{k})}>0$, which further indicates that the parameter estimate $\hat{\theta}_{i}$ is bounded over the interval $t\in[0,\tau_{\textrm{max}})$. Since ${\bm{\varepsilon}}_{1}$ is bounded, and using the properties of $s_{j}$, it is shown that $1\leq-\underline{\delta}_{j}<-\underline{\delta}_{1j}\leq\zeta_{j}(t)\leq\bar{\delta}_{1j}<\bar{\delta}_{j}\leq 1$ for $t\in[0,\tau_{\textrm{max}})$. As $\zeta_{j}=\frac{\eta_{j}}{\varphi_{j}}$ and $0<\varphi_{jf}<\varphi_{j}(t)\leq\varphi_{j0}\leq 1$, it is proved that there exist some constants $\underline{\eta}_{j}$ and $\bar{\eta}_{j}$ so that $1<\underline{\eta}_{j}\leq\eta_{j}\leq\bar{\eta}_{j}<1$, which further implies from $e_{j}=\frac{l_{j}\eta_{j}}{\sqrt{1-\eta_{j}^{2}}}$ that the tracking error $e_{j}$ is bounded for $t\in[0,\tau_{\textrm{max}})$. As $e_{j}=x_{1j}-y_{dj}$, one has $x_{1j}$ is bounded since $y_{dj}$ is bounded for $t\in[0,\tau_{\textrm{max}})$. Hence, it is trivial to prove that $0<\underline{\mu}_{j}\leq\mu_{j}\leq\bar{\mu}_{j}$ and $0<\underline{r}_{j}\leq r_{j}\leq\bar{r}_{j}$, $0<\underline{w}_{j}\leq w_{j}\leq\bar{w}_{j}$ for $t\in[0,\tau_{\textrm{max}})$, with $\underline{\mu}_{j}$, $\bar{\mu}_{j}$, $\underline{r}_{j}$, $\bar{r}_{j}$$\underline{w}_{j}$, $\bar{w}_{j}$ being some positive constants, which indicates that $\bm{W}$ and $\bm{V}$ are bounded. Since $x_{1}$ is bounded, it is shown from Assumption 1 and 4 that $\Phi_{1}$ is bounded, and thus $\bm{a}_{1}$ and $\dot{\hat{\theta}}_{1}$ are bounded by (21) and (22). Following this line of reasoning, $\bm{a}_{i}$ ($i=2,\ldots,N-1$), $\bm{u}$, $\dot{\hat{\theta}}_{k}$ ($k=2,\ldots,N$) and $\bm{x}_{k}$ are bounded on $t\in[0,\tau_{\textrm{max}})$. Importantly, the boundedness of these signals depends solely on the design parameters and is independent of $\tau_{\textrm{max}}$, implying that $\tau_{\textrm{max}}$ can be extended to infinity. Therefore, all signals in the closed-loop system remain bounded for all $t\in[0,\infty)$.
@@
We now demonstrate that the tracking error remains within the prescribed bounds for all $t\in[0,\infty)$. Since $s_{j}$ is bounded over this interval, it follows from Lemma 1 that the tracking error satisfies $\mathcal{H}(-\underline{\delta}_{j}\varphi_{j}(t))<e_{j}(t)<\mathcal{H}(\bar{\delta}_{j}\varphi_{j}(t))$ for all $j=1,2,\ldots,m$. This completes the proof. $\square$
@@
By introducing the unified prescribed performance function $\mathcal{H}$, the originally constrained error dynamics $\dot{\bm{e}}$ are converted into an unconstrained transformed system as shown in (15). The proposed virtual controllers (21), (32), (38) together with the actual controller (46) then stabilize this transformed system, which indirectly enforces the predefined performance constraints on the tracking error. Notably, suitable selection of design parameters enables global or semi-global asymmetric performance to be realized without altering the control framework. Moreover, unexpected actuator faults are compensated automatically without the need for additional fault detection or diagnosis mechanisms.
@@
The controllability condition and controller constructed in this paper can be flexibly applied to a class of square and non-square systems. Specifically, square and non-square systems can be handled by simply choosing $\bm{A}=\bm{I}_{n}$ and $\bm{A}=[\bm{I}_{n},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{(m-n)}]$, respectively. Moreover, since the auxiliary matrix $\bm{P}_{i}$ is adaptively compensated and does not use in the controller, even if $\bm{g}_{i}$ satisfies the original strong controllability condition in [9], it is sufficient to compensate $\bm{P}_{i}$ as a identity matrix $\bm{I}_{n}$ without any additional operation. Notably, the relaxed controllability condition proposed in this work is practically meaningful, as demonstrated in spacecraft applications [23]. Although it does not alter the structure of existing controllers (e.g., [9, 2]), it is clear from the preceding analysis that introducing the unknown auxiliary matrix $\bm{P}_{i}$ can significantly affect both controller design and stability analysis. In this work, such an unknown component is cleverly integrated with system uncertainties, allowing the controller to retain a unified and general structure, while extending applicability to more general and uncertain systems.
@@ V Illustrative Example
V-A Application to quadrotors (square case)
In this section, we consider a quadrotor [17] whose 3-DOF rotational dynamics are described by
$\dot{\bm{\Theta}}=$ $\bm{R}\bm{Q}$ (53)
$\bm{M}\dot{\bm{Q}}=$ $-\bm{Q}\times\bm{M}\bm{Q}+\bm{\tau}+\bm{\Delta}(t)$
where $\bm{\Theta}=[\phi,\theta,\psi]^{T}\in\mathbb{R}^{3}$ represents the Euler angles, $\bm{Q}=[p,q,r]^{T}\in\mathbb{R}^{3}$ denotes the angular velocity, and $\bm{\Delta}(t)\in\mathbb{R}^{3}$ is the disturbance vector. The control torque vector is given by $\bm{\tau}=[\tau_{1},\tau_{2},\tau_{3}]^{T}\in\mathbb{R}^{3}$, the inertia matrix is $\bm{M}=\operatorname{diag}\{M_{xx},M_{yy},M_{zz}\}\in\mathbb{R}^{3\times 3}$, and $\bm{R}\in\mathbb{R}^{3\times 3}$ denotes the transformation matrix. Further details on this model can be found in [3]. Defining $\bm{x}_{1}=\bm{\Theta}$, $\bm{x}_{2}=\dot{\bm{\Theta}}$, and $\bm{X}_{2}=[\bm{x}_{1}^{T},\bm{x}_{2}^{T}]^{T}$, the system (53) can be equivalently rewritten as
$\dot{\bm{x}}_{1}=$ $\bm{x}_{2}$ (54)
$\dot{\bm{x}}_{2}=$ $\bm{f}_{2}(\bm{X}_{2})+\bm{g}_{2}(\bm{x}_{1})\bm{u}_{a}+\bm{d}_{2}(\bm{x}_{1},t)$
with $\bm{f}_{1}(\cdot)=\bm{0}_{3}$, $\bm{f}_{2}(\cdot)=-\bm{R}\bm{M}^{-1}\bm{R}^{-1}\bm{x}_{2}\times\bm{M}\bm{R}^{-1}\bm{x}_{2}+\dot{\bm{R}}\bm{R}^{-1}\bm{x}_{2}$, $\bm{g}_{1}(\cdot)=\bm{I}_{3}$, $\bm{u}_{a}=\bm{\tau}$, $\bm{d}_{1}(\cdot)=\bm{0}_{3}$, $\bm{d}_{2}(\cdot)=\bm{R}\bm{M}^{-1}\bm{\Delta}(t)$, and $\bm{g}_{2}(\cdot)=\bm{R}\bm{M}^{-1}$ calculated as
$\bm{g}_{2}(\bm{x}_{1})=\begin{bmatrix}\frac{1}{M_{xx}}&\frac{\sin(\phi)\tan(\theta)}{M_{yy}}&\frac{\cos(\phi)\tan(\theta)}{M_{zz}}\\ 0&\frac{\cos(\phi)}{M_{yy}}&\frac{-\sin(\phi)}{M_{zz}}\\ 0&\frac{\sin(\phi)\sec(\theta)}{M_{yy}}&\frac{\cos(\phi)\sec(\theta)}{M_{zz}}\end{bmatrix}$. (55)
The fault profiles are outlined as follows:
[…]
The quadrotor’s physical parameters are given by $M_{xx}=0.021\textrm{kg}\cdot\textrm{m}^{2}$, $M_{yy}=0.021\textrm{kg}\cdot\textrm{m}^{2}$, and $M_{zz}=0.039\textrm{kg}\cdot\textrm{m}^{2}$. The desired reference trajectory is $\bm{y}_{d}=[\cos(t),\sin(2t),0.1\tanh(t)]^{T}$, and the disturbance vector is defined as $\bm{\Delta}(t)=[0.02\sin(t),0.02\cos(t),\tanh(t)]^{T}$. For this square system, the actuator allocation matrix is the identity, $\bm{A}=\bm{I}_{n}$. Although the matrix $\bm{\rho}\bm{g}_{2}+\bm{g}_{2}^{T}\bm{\rho}$ is not uniformly positive/negative definite for certain states $\bm{x}_{1}$, the system still meets Assumption 3 by selecting $\bm{P}_{2}=\textrm{diag}\{M_{xx}+\cos(\phi)^{2},M_{yy}+\sec(\theta),M_{zz}+\frac{M_{yy}\cos(\theta)}{M_{zz}\cos(\phi)}\}$.
Figure 2: Control input signals under $\bm{x}_{1}(0)=[0.5,0.5,0.5]^{T}$.
The rate function is chosen as $\beta(t)=\exp(-0.9t)\cos^{2}(t)$, and the key design parameters are selected as $\underline{\delta}_{j}=\bar{\delta}_{j}=1$, $\varphi_{j0}=1$, $\varphi_{jf}=0.1$ and $l=0.9936$, $j=1,2,3$. The other parameters are given as: $\kappa_{1}=1$, $\kappa_{2}=1$, $\sigma_{1}=\sigma_{2}=0.01$ and $\mu_{1}=\mu_{2}=0.1$. The initial conditions are set as: $\bm{x}_{1}(0)=[0.5,0.5,0.5]^{T}$(rad) and $1.2,-0.5,-0.5]^{T}$(rad), $\bm{x}_{2}(0)=[0,0,0]^{T}$, $\hat{\theta}_{1}=\hat{\theta}_{2}=0$. The simulation results are shown in Figs.1 and 2. Fig.1 shows that the output tracking errors $e_{j}$, $j=1,2,3$, evolve within the prescribed performance bounds ($\mathcal{H}(-\varphi_{j}(t)),\mathcal{H}(\varphi_{j}(t))$). Fig.2 shows the boundedness of the control signal $u$.
V-B Application to spacecraft (non-square case)
@@
$\bm{J}_{0}=\begin{bmatrix}20&1.2&0.9\\ 1.2&5&1.4\\ 0.9&1.4&5\end{bmatrix}\textrm{kg}\cdot\textrm{m}^{2}$
@@
[…]
The key design parameters are selected as $\underline{\delta}_{j}=\bar{\delta}_{j}=1$, $j=1,2,3$. The control parameters are given as: $\kappa_{1}=1$, $\sigma_{1}=0.01$ and $\mu_{1}=0.1$. Since configuration matrix $\bm{D}$ can also be interpreted as the actuator allocation matrix, we set $\bm{A}=\bm{D}$. Similarly, it can be verified that $\bm{{g}}\bm{\rho}\bm{D}^{T}+\bm{D}\bm{\rho}\bm{{g}}^{T}$} is not uniformly positive definite or negative definite for all $t\in(0,\infty]$, but we can still make Assumption 3 hold by choosing $\bm{P}=\textmd{diag}\big{\{}0.7,0.1,0.6\big{\}}$.
The rate function is chosen as $\beta(t)=\exp(-0.9t)\cos^{2}(t)$, and $\varphi_{j0}=1$, $\varphi_{jf}=0.1$ and $l=0.9936$, $j=1,2,3$. The initial conditions are set as: $\bm{x}_{1}(0)=[-0.5,-0.5,-0.5]^{T}$(rad) and $1,0.5,0.5]^{T}$(rad), $\hat{\theta}_{1}=0$. The simulation results are given in Figs.3 and 4. Fig.3 shows that the output tracking errors $e_{j}$, $j=1,2,3$, evolve within the prescribed performance bounds ($\mathcal{H}(-\varphi_{j}(t)),\mathcal{H}(\varphi_{j}(t))$). Fig.4 shows the boundedness of the control signal $u$.
To verify that the proposed control is able to achieve the asymmetric performance behaviors, we select the key design parameters as $\underline{\delta}_{j}=0.8,\ \bar{\delta}_{j}=1$ and $\underline{\delta}_{j}=1,\ \bar{\delta}_{j}=0.6$, $j=1,2,3$. The initial conditions are set as: $\bm{x}_{1}(0)=[-0.5,-0.5,-0.5]^{T}$(rad) and $1,0.5,0.5]^{T}$(rad). Other parameters and conditions are kept consistent with the previous simulations. The responses of tracking errors are depicted in Fig.5 and Fig.6, which shows that the tracking errors are always evolve within the prescribed performance bounds, in line with the theoretical analysis.
Figure 6: The evolution of $e_{j}$ under $\underline{\delta}_{j}=1,\ \bar{\delta}_{j}=0.6$, $i=1,2,3$.